Andretti Company has a single product called a Dak. The company normally produces and sells 80,000 Daks each year at a sellin5. An outside manufacturer has offered to produce 80,000 Daks and ship them directly to Andrettis customers. If Andretti ConComplete this question by entering your answers in the tabs below. Req 1A Reg 1B Req 2 Req3 Req 4A to 4C Req 4D Req 5 AssumeComplete this question by entering your answers in the tabs below. Req 1A Reg 1B Reg 2 Req3 Req 4A to 4C Req 4D Req 5 AssumeComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req2 Req 3 Req 4A to 4C Req 4D Reg 5 The comReq 1A Req 1B Req 2 Req 3 Req 4A to 4C Req 4D Req 5 Due to a strike in its suppliers plant, Andretti Company is unable to puComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req2 Req3 Req 4A to 40 Req 4D Req 5 Due to aComplete this question by entering your answers in the tabs below. Req 1A Reg 1B Req 2 Req3 Reg 4A to 4C Req 4D Reg 5 An outs
Answer #1
Contribution margin
selling price per unit 56
less Variable expenses
direct materials 8.5
direct labor 10
Variable manufacturing overhead 2.8
variable selling expense 2.7 24
Contribution margin per unit 32
Req 1A increased sales in units 80000*20%= 16000
contribution margin per unit 32
incremental contribution margin 512000
less added fixed selling expense 150,000
incremental net operarting income 362,000
1-b) Yes
Req 2 Break even price per unit
Variable manufacturing cost per unit 21.3
Shipping cost 1.9
import duties 2.7
permits &licences 0.9
Break even price per unit 26.8 answer
Req 3 Relevant unit cost $2.70 per unit
4) Foregone contribution margin (3333*32) 106656.00
total avoidable fixed cost
fixed manufacturing overhead cost (720000*2/12)*65% 78000
fixed selling cost (320000*2/12)*20% 10667 88666.67
Financial disadvantage -17989.33
80,000*2/12*25%= 3333.333 units
No
5) Variable manfuacturing costs 21.3
fixed manufacturing overhead cost (9*30%)= 2.7
variable selling expense 2.7*1/3 0.90
total costs avoided 24.90